# Imaging two-body correlations in atomic nuclei via low- and high-energy processes

### v1

@@ C.1 Definition
$\mathcal{Q}^{(2)}_{\ell}$ $\equiv(Q^{(1)}_{\ell}\cdot Q^{(1)}_{\ell})$ (48)
$\equiv\sum_{\mu=-\ell}^{\ell}(-1)^{\mu}Q^{(1)}_{\ell\mu}Q^{(1)}_{\ell-\mu}$
@@
C.2 Classical rigid rotor interpretation
@@
Calculating $\mathcal{B}^{2}_{2}(\text{LE})$ from PHFB or the PGCM calculations, the empirical radius $R^{2}_{0}=1.2A^{1/3}$ is replaced by the computed PHFB or PGCM mean-square radius $\langle r^{2}\rangle$. Doing so, the connection between the radius $R_{0}$ of a uniform sphere and the mean-square radius is given by
$\langle r^{2}\rangle=\frac{3}{5}R^{2}_{0}\;.$ (52)
Appendix D Numerical results
@@
D.1 Technical details
The systematic analysis of all even-even nuclei between carbon and nickel is performed within the PHFB approximation based on the deformed HFB minimum and the EM1.8/2.0 chiral Hamiltonian [24]. The three-body interaction is approximated via the rank-reduction method developed in Ref. [12]. A one-body sHO basis is employed, characterized by the frequency $\hbar\omega=12$ MeV, known to be optimal for nuclei in the light to mid-mass regime. All single-particle basis states up to $e_{\mathrm{max}}=10$ are included, while the representation of the initial three-body interaction operator is further restricted to three-body states up to $e_{\mathrm{3max}}=24$.
The nuclei 12C, 16O and 20,22Ne are further studied via PGCM calculations based on axially deformed Bogoliubov states [15, 14, 13]. To achieve an improved treatment of pairing correlations, the reference states entering the PGCM ansatz are generated using variation after particle-number projection, with subsequent projection onto good angular momentum. The single-particle basis is further restricted to $e_{\mathrm{max}}=8$. As demonstrated in Figs. 5 and 6 for PHFB calculations, the reduced two-body mean-square eccentricities $\mathcal{B}^{2}_{2,3}(\mathrm{HE})$ of present interest are well converged at $e_{\mathrm{max}}=8$.
@@ D.1 Technical details
D.2 One-body contribution $\langle\epsilon^{2}_{2}\rangle^{\mathrm{1b}}$
@@
As stipulated by Eq. (44), the classical RR model suggests that the one-body contribution to the normalized mean-squared eccentricity $\langle\epsilon^{2}_{2}\rangle^{\mathrm{1b}}$ is independent of the intrinsic deformation and equal to $1/A$. Figure 7 does display the product $A\langle\epsilon^{2}_{2}\rangle^{\mathrm{1b}}$ from systematic PHFB calculations of nuclear ground states. The product varies only very moderately with $A$, following quite closely the RR approximation.
@@
Figure 7: Ground-state product $\langle\epsilon^{2}_{2}\rangle^{\mathrm{1b}}A$ from systematic PHFB calculations as a function of the mass number.
@@
In Fig. 8 the independence of the product $A\langle\epsilon^{2}_{2}\rangle^{\mathrm{1b}}$ with respect to the intrinsic deformation is further confirmed.
Figure 8: Same as Fig. 7 as a function of the intrinsic deformation parameter $\beta_{20}(\text{dHFB})$.
D.3 Two-body contribution $\langle\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$
As discussed in the body of the paper, the RR model predicts the two-body contribution $\langle\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$ to the normalized mean-square eccentricity to scale with the square of the deformation parameter $\beta_{20}$ (Eq. (44)). This is indeed what is obtained for a quantum rotor, i.e., within the PHFB approximation, as shown in Fig. 9. In addition, a negative offset, not predicted by the classical rotor, is seen to shift that parabola. This offset takes its origin in the Pauli exclusion principle materialized by the exchange term in the two-body density matrix [7] and is indeed absent from the classical RR model used to derive Eq. (44).
@@ D.3 Two-body contribution $\langle\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$
Figure 9: Ground-state value of $\langle\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$ from systematic PHFB calculations as a function of the intrinsic axial quadrupole deformation $\beta_{20}$ of the underlying dHFB state.
Isolating $\langle\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$ for ”spherical” nuclei characterized by $\beta_{20}^{2}(\text{dHFB})\approx 0$, Fig. 10 shows that the magnitude of the negative offset decreases with $A$. This particular property needs to be confirmed analytically [6].
@@
Figure 10: Ground-state value of $\langle\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$ from systematic PHFB calculations as a function of AA for nearly ”spherical” nuclei characterized by $\beta_{20}^{2}(\text{dHFB})\approx 0$.
@@
A hallmark of a rigid rotor is the characteristic energy ratio $R_{4/2}\equiv E^{*}(4^{+}_{1})/E^{*}(2^{+}_{1})=10/3$, with $E^{*}$ denoting the excitation energy. Figure 11 displays the $R_{4/2}$ ratio from our systematic PHFB calculations as a function of $\mathcal{B}^{2}_{2}(\mathrm{HE})$. For nuclei with moderate to large values of $\mathcal{B}^{2}_{2}(\mathrm{HE})$, the rigid rotor picture is indeed valid. Contrarily, spherical systems identified by their negative $\mathcal{B}^{2}_{2}(\mathrm{HE})$ value show the largest deviation from the pure rigid rotor behavior.

### v2

@@ III.1 Definition
$\mathcal{Q}^{(2)}_{\ell}$ $\equiv(Q^{(1)}_{\ell}\cdot Q^{(1)}_{\ell})$
$\equiv\sum_{\mu=-\ell}^{\ell}(-1)^{\mu}Q^{(1)}_{\ell\mu}Q^{(1)}_{\ell-\mu}$ (57)
@@
III.2 Classical rigid rotor interpretation
@@
$\mathcal{B}^{2}_{2}(\text{LE})$ $\equiv\frac{16\pi^{2}}{9Z^{2}e^{2}R_{0}^{4}}\langle\mathcal{Q}^{(2)}_{2}\rangle$
$=\frac{16\pi^{2}A^{2}}{25Z^{2}e^{2}}\frac{\langle\mathcal{Q}^{(2)}_{2}\rangle}{\langle r^{2}\rangle^{2}}\;,$ (60)
where the empirical radius $R^{2}_{0}=1.2A^{1/3}$ has been replaced by the computed PHFB or PGCM (non-normalized [^6] Mass mean-squared radii compared to experimental data include a normalization; e.g., the proton number for the point proton mean-square radius or the mass number for the mass mean-square radius. No such normalization is used in the present study when defining the expectation value of a squared radius operator.) mean-square radius $\langle r^{2}\rangle$. Doing so, the connection between the radius $R_{0}$ of a uniform sphere and the mean-square radius is given by
$\langle r^{2}\rangle=\frac{3}{5}AR^{2}_{0}\;.$ (61)
IV Numerical results
@@
IV.1 Technical details
The systematic analysis of even-even nuclei between carbon and nickel is performed within the PHFB approximation based on the deformed HFB minimum and the EM1.8/2.0 chiral Hamiltonian [28]. The three-body interaction is approximated via the rank-reduction method developed in Ref. [16]. A one-body sHO basis is employed, characterized by the frequency $\hbar\omega=12$ MeV, known to be optimal for nuclei in the light to mid-mass regime. All single-particle basis states up to $e_{\mathrm{max}}=10$ are included, while the representation of the initial three-body interaction operator is further restricted to three-body states up to $e_{\mathrm{3max}}=24$.
The nuclei 12C, 16O and 20,22Ne are further studied via PGCM calculations based on axially deformed Bogoliubov states [19, 18, 17]. To achieve an improved treatment of pairing correlations, the reference states entering the PGCM ansatz are generated using variation after particle-number projection, with subsequent projection onto good angular momentum and parity. The single-particle basis is further restricted to $e_{\mathrm{max}}=8$. As demonstrated in Figs. 5 and 6 for PHFB calculations, the rescaled two-body mean-square eccentricities $\mathcal{B}^{2}_{2,3}(\mathrm{HE})$ of present interest are well converged at $e_{\mathrm{max}}=8$.
@@ IV.1 Technical details
IV.2 One-body contribution $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{1b}}$
@@
As stipulated by Eq. (53), the classical RR model suggests that the one-body contribution to the normalized mean-squared eccentricity $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{1b}}$ is independent of the intrinsic deformation and equal to $1/A$. Figure 7 does display the product $A\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{1b}}$ from systematic PHFB calculations of nuclear ground states. The product varies only very moderately with $A$, following quite closely the RR approximation. Clearly, the few PGCM points visible on the figure remain consistent with this observation.
@@
Figure 7: Ground-state product $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{1b}}A$ from PHFB and PGCM calculations as a function of the mass number.
@@
In Fig. 8, the independence of the product $A\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{1b}}$ with respect to the intrinsic deformation is further confirmed. One observes that the shape fluctuations included in the PGCM wave-function ansatz do not modify this property.
@@
Figure 8: Same as Fig. 7 as a function of the intrinsic axial quadrupole deformation $\beta_{20}$ of the underlying dHFB state. PGCM results are plotted using the same $\beta_{20}$ values as the PHFB calculations.
IV.3 Two-body contribution $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$
As discussed in the body of the paper, the RR model predicts the two-body contribution $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$ to the normalized mean-square eccentricity to scale with the square of the deformation parameter $\beta_{20}$ (Eq. (53)). This is indeed what is obtained for a quantum rotor, i.e., within the PHFB approximation, as shown in Fig. 9. In addition, a negative offset, not predicted by the classical rotor, is seen to shift that parabola. This offset takes its origin in the Pauli exclusion principle materialized by the exchange term in the two-body density matrix [8] and is indeed absent from the classical RR model used to derive Eq. (53). Because the (quantum) rotor character of the many-body state is altered by shape fluctuations, the PGCM results depart slightly from the PHFB trend.
@@ IV.3 Two-body contribution $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$
Figure 9: Ground-state value of $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$ from PHFB and PGCM calculations as a function of the intrinsic axial quadrupole deformation $\beta_{20}$ of the underlying dHFB state. PGCM results are plotted using the same $\beta_{20}$ values as the PHFB calculations.
Isolating $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$ for “spherical” nuclei characterized by $\beta_{20}^{2}(\text{dHFB})\approx 0$, Fig. 10 shows that the magnitude of the negative offset decreases with $A$. This particular property needs to be confirmed analytically [7].
Figure 10: Ground-state value of $\langle\delta\epsilon^{2}_{2}\rangle^{\mathrm{2b}}$ from PHFB and PGCM calculations as a function of AA for nearly “spherical” nuclei characterized by $\beta_{20}^{2}(\text{dHFB})\approx 0$.
@@
A hallmark of a rigid rotor is the characteristic energy ratio $R_{4/2}\equiv E^{*}(4^{+}_{1})/E^{*}(2^{+}_{1})=10/3$, with $E^{*}$ denoting the excitation energy. Figure 11 displays the $R_{4/2}$ ratio from our PHFB and PGCM calculations as a function of $\mathcal{B}^{2}_{2}(\mathrm{HE})$. For nuclei with moderate to large values of $\mathcal{B}^{2}_{2}(\mathrm{HE})$, the rigid rotor picture is indeed valid. Contrarily, spherical systems identified by their negative $\mathcal{B}^{2}_{2}(\mathrm{HE})$ value show the largest deviation from the pure rigid rotor behavior.
@@ IV.4 Excited $0^{+}_{2}$ state in 12C, 16O and 20,22Ne
IV.5 Kumar effective deformation parameter
As stipulated in the main-body of the paper, the inclusion of the trivial contribution originating from the one-body part of the Kumar operator $\mathcal{Q}^{(2)}_{\ell}$ into the definition of the effective deformation parameter $\mathcal{B}^{2}_{\ell}(\mathrm{LE})$ is responsible for the fact that the latter does not correlate well with the intrinsic deformation parameter $\beta^{2}_{\ell 0}$ of the (quantum) rotor model, i.e., the PHFB approximation. This is at variance with the effective deformation parameter $\mathcal{B}^{2}_{\ell}(\mathrm{HE})$ that does not include such a one-body term and that correlates well with $\beta^{2}_{\ell 0}$. The marked difference between both effective deformation parameters demonstrate that the inclusion of the completeness relation to express $\langle\mathcal{Q}^{(2)}_{\ell}\rangle$ in terms of reduced electromagnetic transitions probabilities and the application of the RR model to infer an effective deformation parameter from it constitute two operations that do not commute. Indeed, applying the RR model after the inclusion of the completeness relation wrongly delivers $\langle\mathcal{Q}^{(2)}_{\ell}\rangle^{\text{RR}}_{0^{+}}\propto\beta^{2}_{\ell 0}$ and thus leads to overlooking the contribution of the one-body term to $\langle\mathcal{Q}^{(2)}_{\ell}\rangle_{0^{+}}$.
This key difference between both effective deformation parameters is presently investigated in more detail, both analytically and numerically. For this purpose, another effective parameter $\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})$ is defined by wrongly including the one-body contribution to the mean-square eccentricity according to
$\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})$ $\equiv\frac{4\pi}{3}\langle\delta\epsilon^{2}_{2}\rangle$
$=\frac{2\pi}{3}\frac{\langle[\mathcal{E}^{(2)}_{2}]_{00}\rangle}{\langle R^{(1)}_{2}\rangle^{2}}$
$=\frac{16\pi^{2}}{25}\frac{\langle\mathcal{Q}^{(2)}_{2}(p+n)\rangle}{\langle r^{2}\rangle^{2}}\;,$ (62)
where $\mathcal{Q}^{(2)}_{2}(p+n)$ extends the definition of the Kumar operator $\mathcal{Q}^{(2)}_{2}$ by summing over both protons and neutrons.
To quantify the role of the incriminated one-body term, two different ratios of effective deformation parameters
$\frac{\mathcal{B}^{2}_{2}(\mathrm{LE})}{\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})}$ $=\frac{4\langle\mathcal{Q}^{(2)}_{2}\rangle}{\langle\mathcal{Q}^{(2)}_{2}(p+n)\rangle}\,$, (63a)
$\frac{\mathcal{B}^{2}_{2}(\mathrm{LE})}{\mathcal{B}^{2}_{2}(\mathrm{HE})}$ $=\frac{4\langle\mathcal{Q}^{(2)}_{2}\rangle}{\langle\mathcal{Q}^{(2)\,\mathrm{2b}}_{2}(p+n)\rangle}\,$, (63b)
are investigated. To mitigate the fact that $\mathcal{B}^{2}_{2}(\mathrm{LE})$ involves only protons while $\mathcal{B}^{2}_{2}(\mathrm{HE})$ and $\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})$ sum over both protons and neutrons, the analysis is presently restricted to $Z=N$ nuclei. In this case, the one-body and two-body components of the operator $\mathcal{Q}^{(2)}_{2}(p+n)$ satisfy [^7] These identities are strictly valid if omitting the Coulomb interaction and the isospin breaking contributions to the strong interaction.
$\mathcal{Q}^{(2)\,\mathrm{1b}}_{2}(p+n)=2\mathcal{Q}^{(2)\,\mathrm{1b}}_{2}\;,$ (64a)
$\mathcal{Q}^{(2)\,\mathrm{2b}}_{2}(p+n)=4\mathcal{Q}^{(2)\,\mathrm{2b}}_{2}\;.$ (64b)
such that both ratios simplify to
$\left[\frac{\mathcal{B}^{2}_{2}(\mathrm{LE})}{\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})}\right]_{N=Z}$ $=\frac{4\langle\mathcal{Q}^{(2)\,\mathrm{1b}}_{2}\rangle+4\langle\mathcal{Q}^{(2)\,\mathrm{2b}}_{2}\rangle}{2\langle\mathcal{Q}^{(2)\,\mathrm{1b}}_{2}\rangle+4\langle\mathcal{Q}^{(2)\,\mathrm{2b}}_{2}\rangle}\;,$ (65a)
$\left[\frac{\mathcal{B}^{2}_{2}(\mathrm{LE})}{\mathcal{B}^{2}_{2}(\mathrm{HE})}\right]_{N=Z}$ $=\frac{4\langle\mathcal{Q}^{(2)\,\mathrm{1b}}_{2}\rangle+4\langle\mathcal{Q}^{(2)\,\mathrm{2b}}_{2}\rangle}{4\langle\mathcal{Q}^{(2)\,\mathrm{2b}}_{2}\rangle}\,.$ (65b)
Further applying the RR model eventually leads to the characteristic forms
$\left[\frac{\mathcal{B}^{2}_{2}(\mathrm{LE})}{\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})}\right]^{\mathrm{RR}}_{N=Z}$ $=1+\frac{1}{1+\frac{3A\beta^{2}_{20}}{8\pi}}\,$, (66a)
$\left[\frac{\mathcal{B}^{2}_{2}(\mathrm{LE})}{\mathcal{B}^{2}_{2}(\mathrm{HE})}\right]^{\mathrm{RR}}_{N=Z}$ $=1+\frac{8\pi}{3A\beta^{2}_{20}}\,.$ (66b)
Both ratios are displayed in Fig. 12 for PHFB results of $N=Z$ nuclei between carbon and nickel as a function of the product $A\cdot\beta^{2}_{20}(\text{dHFB})$. The observed trend of both ratios is fully consistent with the RR analytical result obtained in Eq. (66).
Figure 12: Ratios $\mathcal{B}^{2}_{2}(\mathrm{LE})/\mathcal{B}^{2}_{2}(\mathrm{HE})$ and $\mathcal{B}^{2}_{2}(\mathrm{LE})/\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})$ from PHFB calculations of $N=Z$ nuclei between carbon and nickel as a function of $A\cdot\beta^{2}_{20}(\mathrm{dHFB})$. The inset zooms over the ratios ranging between $1.0$ and $2.5$.
Overall, the $\mathcal{B}^{2}_{2}(\mathrm{LE})/\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})$ is much closer to $1$ than $\mathcal{B}^{2}_{2}(\mathrm{LE})/\mathcal{B}^{2}_{2}(\mathrm{HE})$, which indeed demonstrates that the inappropriate inclusion of the one-body term into $\mathcal{B}^{2}_{2}(\mathrm{LE})$ generates the difference from $\mathcal{B}^{2}_{2}(\mathrm{HE})$. Let us further analyze the behavior of both ratios.
For light nuclei displaying small intrinsic deformation, the dominance of the one-body term combined with the different isospin character of both operators at play result into $\mathcal{B}^{2}_{2}(\mathrm{LE})/\tilde{\mathcal{B}}^{2}_{2}(\mathrm{HE})\sim 2$ (Eq. (66a)) [^8] The negative contribution of the exchange term in spherical systems generates in fact values that are slightly larger than 2.. As predicted by the RR model (Eq. (66b)), the ratio decreases towards 1 as the product $A\cdot\beta^{2}_{20}(\mathrm{dHFB})$ increases due to the dominance of the two-body contribution to both operators in this limit. However, in the set of nuclei under study there is no system for which the two-body term is sufficiently large to effectively approach 1; i.e., for the largest value of $A\cdot\beta^{2}_{20}(\mathrm{dHFB})$ presently reached, the ratio remains half way between $2$ and $1$ as the inset shows.
More importantly, the relevant ratio $\mathcal{B}^{2}_{2}(\mathrm{LE})/\mathcal{B}^{2}_{2}(\mathrm{HE})$ is much larger than $1$ for most of the nuclei under study, thus confirming the inoperative character of the effective deformation parameter $\mathcal{B}^{2}_{2}(\mathrm{LE})$ extracted from the Kumar operator. Whenever $\beta^{2}_{20}(\mathrm{dHFB})$ is small, the second term in Eq. (66b) is further enhanced due to the presently neglected exchange term that effectively reduces the denominator [^9] In fact, $\mathcal{B}^{2}_{2}(\mathrm{HE})$ is eventually negative in the limit $\beta^{2}_{20}(\mathrm{HFB})\approx 0$. The corresponding points are not depicted in Fig. 12.. As the product $A\cdot\beta^{2}_{20}(\mathrm{dHFB})$ increases, the ratio $\mathcal{B}^{2}_{2}(\mathrm{LE})/\mathcal{B}^{2}_{2}(\mathrm{HE})$ decreases as $(A\cdot\beta^{2}_{20}(\mathrm{dHFB}))^{-1}$ to eventually reach $1$ in well-deformed heavy nuclei. For the nuclei under study whose maximum mass is $A=80$, this ratio is however never smaller than $3$.